Use the given conditions to write an equation for the line in point-slope form and slope-intercept form.
x intercept= -1/9
y intercept= 7
note- *simplify any fraction as much as possible and do not use decimals*

Found 2 solutions by solver91311, DrBeeee:
Answer by solver91311(24713)



The conditions for this problem are improperly stated in my view. The intercepts of a line are points which are properly described with ordered pairs (assuming that you are working in space) rather than scalar constants.

Hence, the conditions of the problem should have been -intercept: and -intercept:

Use the two-point form of an equation of a line:



where and are the coordinates of the given points.

Do the arithmetic and simplify to the point-slope form:



and the slope-intercept form:



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by DrBeeee(684)

Haven't heard of the point-slope form, but I'll give you the three basic forms:
I Slope-Intercept Form is y = m*x + b, where m is the slope and b is the y-intercept
II Standard Form is A*x + B*y = C, where A, B, and C are constants and
III Intercept Form is x/(x-int) + y/(y-int) = 1, where x-int is the x intercept and y-int is the y intercept.
The third one is the easiest for you to start with because the x-int and y-int are given. Substituting them into III yields
(1) x/(-1/9) + y/(7) = 1
Simplifying (1) gives us
(2) -9*x + y/7 = 1 or
(3) -63*x + y = 7, which is the standard form II.
Solving (3) for y yields
(4) y = 63*x + 7, which is the slope intercept form.
Take your pick!